Question: The asset beta of a levered firm is 1.4. The beta of debt is 0.5. If the debt to value ratio is 0.3, what is the equity beta?

Formula: Equity beta = Asset beta + Debt to equity ratio * (Asset beta - Beta of Debt)

*the debt to equity ratio is NOT given. I must find the debt to equity ratio first to complete the answer. Only the debt to value ratio is provided.
Answer #1

D/E = D/(V-D) = 0.3/(1-0.3)=0.428571

Equity beta = Asset beta + Debt to equity ratio * (Asset beta - Beta of Debt)

=1.4+0.428571*(1.4-0.5)=1.78571